OpenStudy (anonymous):

In addition, it is sometimes possible to monitor certain characteristics of the animals. The length of an alligator can be estimated accurately from aerial photographs or from a boat. However, the alligators weight is much more difficult to determine. The example below provides data on the length (inches) and weight (pounds) of alligators captured in Florida. Develop a model in which the weight of an alligator can be predicted from its length. Show all of your thinking.

OpenStudy (anonymous):

Length Weight 58 28 61 44 63 33 68 39 69 36 72 38 72 61 74 54 74 51 76 42 78 57 82 80 85 84 86 83 86 80 86 90 88 70 89 84 90 106 90 102 94 110 94 130 114 197 128 366 147 640

OpenStudy (anonymous):

Can anyone help out? Because some assistance would really be appreciated.

OpenStudy (anonymous):

@Hero @amistre64 @ybarrap Anyone? Please?

OpenStudy (anonymous):

plot all your data points, then estimate a line

OpenStudy (anonymous):

or search for a linear regression calculator

OpenStudy (anonymous):

What does it mean by develop a model? What am I supposed to do for that? It says to show all my work but how and what?

OpenStudy (anonymous):

Can anyone assist? I desperately need help and soon. This assignment does have a due date and I really need help in order to finish.

OpenStudy (anonymous):

develop a model means find an equation

OpenStudy (amistre64):

excel has a trendline feature my ti83 has a linear regression function as well as some fancier regression functions. if this is a stats class ... you most likely have a program available to you to work with the data set. Statcrunch maybe?

OpenStudy (anonymous):

Nope, no mention of any program. My course doesn't explain a lot of things or offer much assistance often enough. Find an equation - what equation? How? I'm sorry, I'm rather dumb and I'm confused.

OpenStudy (amistre64):

what sort of programs do you have at your disposal? excel perhaps? do you have a fancy-schmancy graphing calculator? because doing this by hand would take about 3 weeks to complete

OpenStudy (amistre64):

we could develop a linear equation which may or may not best fit the data but is the "simplest" to create. It would require knowing: the mean length, the mean weight, the sum of the lengths, the sum of the weights, the sum of the square of the weights, and the sum of the products of length and weight.

OpenStudy (amistre64):

once we have all that information we can construct a linear equation as:\[\hat y=\frac{n\sum xy-\sum x\sum y}{n\sum xx-\sum x\sum x}(x-\mu_x)+\mu_y\]

OpenStudy (anonymous):

I have a graphing calculator.

OpenStudy (anonymous):

I'm sorry, all those things just went right over my head. What?

OpenStudy (amistre64):

see why it would take 3 weeks lol is your calculator a ti83 or similar? i have a ti83

OpenStudy (amistre64):

if you have a ti30 then the process is tedious, but then the ti30 isnt a graphing calc so i dont know why i would even bring that up

OpenStudy (anonymous):

I have a casio fx-9750gII graphing calculator

OpenStudy (amistre64):

bummer ... i got know idea how that one would work :/ do you have lists that you can fill in with it?

OpenStudy (amistre64):

http://www.youtube.com/watch?v=KPEEVdkA7Wk this might be useful to you

OpenStudy (anonymous):

what kind of lists? like entering numbers in the kind of list that you make into a graph kind of list?

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

menu, stat option, ... looks like the lists just pop up for you after that just fill them in with the dataset they provided

OpenStudy (anonymous):

I did that. now what?

OpenStudy (amistre64):

you tube says to select a graph, and itll scatterplot it for you; then calc gets you to the regression options

OpenStudy (amistre64):

when i use excel i get this: the scatter plot looks like an x^2 is a best fit model.

OpenStudy (anonymous):

That's what I got on my calculator. It curves up like that and everything.

OpenStudy (amistre64):

the exponential is a pretty nice fit too, does it ask you for a specific type of curve to model? line, polynomial, exponential, etc ? or are you left to your own judgement for that?

OpenStudy (amistre64):

id go with the poly x^2 or the exponential ...

OpenStudy (anonymous):

All it says is this: Develop a model in which the weight of an alligator can be predicted from its length. Show all of your thinking.

OpenStudy (anonymous):

So, I am confused. How am I supposed to like say my answer? The whole develop a model thing? Do I like say that the exponential is a good fit for the model and that's my answer? or...?

OpenStudy (anonymous):

Should I also include the formula thing for the x^2 curve thing which is y=ax^2+bx+c ? (I think that's what it is anyway. That's right isn't it?)

OpenStudy (anonymous):

@JJ_Stark do u take chemistry?

OpenStudy (anonymous):

Yes @nannaberry

OpenStudy (anonymous):

can i tag u?

OpenStudy (amistre64):

i would say something like, the scatterplot suggests that a linear correlation is not applicable, and that it suggests a quadratic relationship. In my opinion, a quadratic function model best fits the data for predictive purposes.

OpenStudy (anonymous):

Thank you so much. You have been so incredibly helpful(:

OpenStudy (anonymous):

@JJ_Stark would u mind comin to my page now :)